Ch 6 Ch61 17 Mauro Products distributes a single product, a woven basket whose selling price is $12 per unit and whose variable expense is $8 per unit. The companys monthly foxed expense is $10,400 te the companys break-even point in unit sales. 2. Calculate the companys break-even point in dollar sales. (Do not round intermediate calculations.) p3. If the companys fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do n und <Prew 17 of 23Next>
Answer #1

Answer:

1) Profit = unit contribution * quantity - fixed expenses

=> 0 = (12-8) * Q - 10400 (i.e, the profit is 0 at break even point)

=> 0 = 4Q - 10400 => 4Q = 10400 => Q = 2600

Break even point in unit sales = 2600 baskets

2) Dollar Sales is = break-even sales units * sale price

= 2600 baskets * $12 = $31200

break even point in dollar sales is $31200.

3) revised fixed cost = $10400 + $600 = $11000

unit sales = fixed cost / contribution margin = 11000 / 12- 8 = 11000 / 4 = 2750 baskets

break even point in unit sales is 2750 baskets

=> dollar sales = 2750 baskets * $12 = $33000

break even point in dollar sales is $33000.

Break-even point in unit sales Break-even point in dollar sales Break-even point in unit sales Break-even point in dollar sales 1 baskets 2600 $31,200 2750 $33,000 2 baskets